How many moles of silver atoms are in 1.8 times 10 raised to the 20 power atoms of silver
1
Expert's answer
2012-10-12T10:48:39-0400
n = N/NA
calculations:
N - 1.8 * 10^20 atoms of silver
NA = 6.02 * 10^23 mol−1
n = 1.8 * 10^20/ 6.02 * 10^23 = 2.99x 10^-4 moles
answer:
n = 2.99x 10^-4 moles (0,000299 moles)